OpenStudy (anonymous):

Simplify the given expression below: 4 divided by the quantity of 3 minus 2i

OpenStudy (anonymous):

\[\frac{ 4 }{ 3-2i }\]

OpenStudy (anonymous):

@The_One_And_Only wouldn't I multiply the numerator and denominator by -1?

OpenStudy (anonymous):

meant @TheSmartOne

TheSmartOne (thesmartone):

I would think that you would need to multiply the numerator and denominator by the conjugate of 3 - 2i so that way there is no i in the denominator.

OpenStudy (anonymous):

First fiind the complex conjugate of 3-2i and multiply both nummerator and denominator by it:\[\frac{4}{3-2i} \times\frac{conj}{conj}\]

OpenStudy (anonymous):

did that and got \[\frac{ 12-8i }{ 9+2i }\]

TheSmartOne (thesmartone):

\(\sf (a-bi)(a+bi) = a^2 + b^2\) Try again.

TheSmartOne (thesmartone):

The i should be complete gone because i^2 = -1 and (a-b)(a+b) = a^2 - b^2 and then the i^2 is -1 so it makes it a^2 + b^2 as you can see in the formula I gave you

OpenStudy (anonymous):

Not understanding the options give a 13 at the denominator not sure how I would get 13 for it keep getting 7 or 11

TheSmartOne (thesmartone):

\(\sf\Large \frac{4}{3-2i} \times \frac{3+2i}{3+2i} = \frac{4(3+2i)}{(3-2i)(3+2i)}\) Now, \(\sf (a-bi)(a+bi) =a^2 + b^2\\\\(3-2i)(3+2i) = ??\)

OpenStudy (anonymous):

9+6i-6i-4i 9+4i

OpenStudy (anonymous):

would I be canceled out since (a+b)^2

OpenStudy (anonymous):

or put it up with the numerator

TheSmartOne (thesmartone):

\(\ 2i \times 2i = 4i^2\) and \(\sf i^2 = -1\) so \(\sf 4i^2 = -4\)

TheSmartOne (thesmartone):

9 - 4i^2 = 9 - (-4) = 9 + 4 = 13 so now we have: \(\sf\Large \frac{4}{3-2i} \times \frac{3+2i}{3+2i} = \frac{4(3+2i)}{(3-2i)(3+2i)}=\frac{4(3+2i)}{13}\)

OpenStudy (anonymous):

so\[\frac{ 12+8i }{ 13 }\]

OpenStudy (anonymous):

thank yall for the help

TheSmartOne (thesmartone):

Anytime!